Complete the below table. Consumption function or saving function before tax Induced tax Consumption function or saving function after tax C = 1800 + 0.7Yd T = 0.2Y (a) [2m] S = -105+ 0.4Yd T = 0.4Y (b)[2m]


1
Expert's answer
2021-05-25T11:10:23-0400

Solution:

a.). Consumption function before tax:

We first eliminate Taxes from Yd:

Yd = Y – T = Y

C = 1800 + 0.7Y

We know:

Y = C + I + G

Y = C

Y = 1800 + 0.7Y

Y – 0.7Y = 1800

0.3Y = 1800

Y = 6,000

Substitute in the C function to derive Consumption before taxes:

C = 1800 + 0.7(6000)

C = 1800 + 4200

C = 6,000

Consumption function before taxes = 6,000

Consumption function after tax:

C = 1800 + 0.7Yd

Yd = Y – T

T = 0.2Y

C = 1800 + 0.7 (Y – 0.2Y)

Y = C

Y = 1800 + 0.7 (Y – 0.2Y)

Y = 1800 + 0.56Y

Y – 0.56Y = 1800

0.44Y = 1800

Y = 4,091

Substitute in the C function to derive Consumption after taxes:

C = 1800 + 0.7(4091)

C = 1800 + 2864

C = 4,664

Consumption function after taxes = 4,664


b). Saving function before tax:

S = -105 + 0.44Yd

We first eliminate Taxes from Yd:

Yd = Y – T = Y

S = -105 + 0.44Y

We know Y before taxes = 6000:

Substitute Y in the Saving function:

S = -105 + 0.44(6000)

S = -105 + 2400

S = 2,295

Saving function before taxes = 2,295

Saving function after tax:

S = -105 + 0.44Yd

S = -105 + 0.44 (Y – T)

S = -105 + 0.44 (Y – 0.4Y)

S = -105 + 0.24Y

S = -105 + 0.24(4091)

S = -105 + 982

S = 877

Saving function after taxes = 877